OpenStudy (anonymous):

find an equation of the line that passes through the point (-3, -2) and is parallel to 2x = 5y

OpenStudy (anonymous):

y = 2/5 x + b

OpenStudy (anonymous):

To determine the slope of 5y = 2x, solve for y. You get y = 2/5 X. Now, substitute the point & slope into y = mx + b. solve for b. b = -2 2/5. Your equation for the line is y = -2/5 x - 2 2/5